assume that when adults with smartphones are randomly selected, 56% use them in meetings or classes. If 12 adult smartphone users are randomly selected, find the probability that fewer that 5 of them use their smartphones in meetings or classes
Answer #1

Solution :

Given that n = 12 , p = 0.56

=> q = 1 - p = 0.44

=> For binomial distribution, P(x,r) = nCr*p^r*q^(n - r)

=> P(x < 5) = P(x <= 4)

= P(x = 0) + P(x = 1) + P(x = 2) + P(x = 3) + P(x = 4)

= 12C0*0.56^0*0.44^12 + 12C1*0.56^1*0.44^11 + 12C2*0.56^2*0.44^10 + 12C3*0.56^3*0.44^9 + 12C4*0.56^4*0.44^8

= 0.0001 + 0.0008 + 0.0056 + 0.0239 + 0.0684

= 0.0988